OpenStudy (anonymous):

Hi im working in math and have a hard time working through math questions can someone please EXPLAIN how to do this question please? A new movie theater has stadium seating. The first row has 10 seats. Each additional row has 4 seats more than the row before it. How many seats are in the theater if there are 16 rows of seats? Thx

OpenStudy (amistre64):

this is representative of an arithmetic sequence .... a list of numbers that differ by some constant adder

OpenStudy (amistre64):

row1 = 10 row2 = 10+4 row3 = 10 + 4 + 4 etc ...

OpenStudy (amistre64):

if we notice a pattern in the 4s, there are n-1 4s in the nth row row4 therefore has 10 + 4(3) seats row5 has 10+ 4(4) seats and so on

OpenStudy (amistre64):

oh, and if i read all the way thru it .... lol add up all the rows to find the total number of seats

OpenStudy (amistre64):

so, 10 + 4( 0) + 10 + 4( 1) + 10 + 4( 2) + 10 + 4( 3) + 10 + 4( 4) ... + 10 + 4( 16) ------------- 17(10) + 4(0+1+2+3+4+...+16)

OpenStudy (anonymous):

so would it be 72